# Multiconnectivity for SAGIN: Current Trends, Challenges, AI-driven Solutions, and Opportunities

### v1

@@ Abstract
Space-air-ground-integrated network (SAGIN)-enabled multiconnectivity (MC) is emerging as a key enabler for next-generation networks, enabling users to simultaneously utilize multiple links across multi-layer non-terrestrial networks (NTN) and multi-radio access technology (multi-RAT) terrestrial networks (TN). However, the heterogeneity of TN and NTN introduces complex architectural challenges that complicate MC implementation. Specifically, the diversity of link types, spanning air-to-air, air-to-space, space-to-space, space-to-ground, and ground-to-ground communications, renders optimal resource allocation highly complex. Recent advancements in reinforcement learning (RL) and agentic artificial intelligence (AI) have shown remarkable effectiveness in optimal decision-making in complex and dynamic environments. In this paper, we review the current developments in SAGIN-enabled MC and outline the key challenges associated with its implementation. We further highlight the transformative potential of AI-driven approaches for resource optimization in a heterogeneous SAGIN environment. To this end, we present a case study on resource allocation optimization enabled by agentic RL for SAGIN-enabled MC involving diverse radio access technologies (RATs). Results show that learning-based methods can effectively handle complex scenarios and substantially enhance network performance in terms of latency and capacity while incurring a moderate increase in power consumption as an acceptable tradeoff. Finally, open research problems and future directions are presented to realize efficient SAGIN-enabled MC.
@@ I-A MC in the Unified and Standardized SAGIN Architecture
Nonetheless, the integration of multi-layer NTNs and multi-RAT TNs introduces complex architectures and a plethora of use cases. Consequently, a variety of link types include air-to-air, air-to-space, space-to-space, space-to-ground, and ground-to-ground links. Fig. 1 illustrates a scenario of the heterogeneous environment and architectures enabled by unified SAGIN, where the coexistence of space, air, and ground platforms is considered. As shown in the figure, it encompasses unmanned aerial vehicles (UAVs), high-altitude platforms (HAPs), satellites, and hybrid gateway-base stations (GS-BSs) with highly dynamic and complex network conditions, which makes it more challenging to optimize resource allocation and management. For instance, rapidly changing channel conditions, frequent handovers across space, air, and ground domains, and varying latency, bandwidth, and reliability requirements are the factors that contribute to the resource management complexity.
@@
  • The proposed algorithm is validated through simulations and compared with relevant baselines. The results demonstrate considerable performance improvements of the considered KPIs and provide insight into the tradeoffs among them. In addition, we present a comprehensive overview of open challenges and potential research directions.
@@ III Agentic RL for SAGIN-enabled MC
Recently, RL-driven agentic AI, called agentic RL has emerged as a key tool for addressing complex network optimization challenges [11]. In the context of SAGIN-enabled MC, such agents optimize resource allocation under dynamic and complex conditions. The agents continuously learn rapidly changing network states, diverse user requirements, and complex mobility patterns. This enables them to adapt to environmental dynamics and facilitate efficient service delivery. The agents also enable proactive decision-making regarding link activation, handover, traffic steering, and load balancing to predict network conditions. This allows for the intelligent definition of dynamic resource allocation patterns and the optimization of parameters such as transmit power, bandwidth, and carrier selection to maximize spectral and energy efficiency. Thus, agentic RL can enable intelligent, self-optimizing control policies for SAGIN-enabled MC. These policies can orchestrate handovers, link selection, resource allocation, and load balancing across heterogeneous domains while ensuring scalability and adaptability under highly nonlinear dynamics.
@@
Consequently, we have $\{\operatorname{\mathcal{S}},\operatorname{\mathcal{A}},p(s^{\prime}|s_{t},a_{t}),r_{t}\}$, where $\operatorname{\mathcal{S}}$ denotes the state space of the environment, while $\operatorname{\mathcal{A}}$ represents the action space for making the optimal decision. In addition, $p(s^{\prime}|s_{t},a_{t})$ corresponds to the probability of transitioning to state $s^{\prime}\in\operatorname{\mathcal{S}}$ conditioned on taking action $a_{t}\in\operatorname{\mathcal{A}}$ in state $s_{t}\in\operatorname{\mathcal{S}}$. The term $r_{t}$ represents the immediate reward obtained by taking action $a_{t}$, which drives the transition from $s_{t}$ to $s^{\prime}$. We assume that a UE has access to $N$ available links, indexed by $n\in\{1,...,N\}$, from which a subset can be dynamically assigned to the UE.
@@ III-C Reward
The immediate reward $r_{t}$ is formulated as a penalty imposed on the agent that is proportional to the capacity, latency, and power usage calculated over an episode. More precisely, the reward at each time step is defined as a weighted combination that increases with higher capacity and decreases with higher latency and power consumption. Thus, the reward function maximizes the achievable transmission rate while minimizing normalized system delay. The function implicitly penalizes packet loss, high delays, and exceeding network load constraints, thereby guiding the agent to learn optimal link selection policies. Furthermore, it encourages selecting fewer links to minimize cost.
@@ III-D Agent Architecture
As illustrated in Fig. 3, the architecture of the proposed learning agent model comprises three neural networks: the actor $A(s_{t})$, the critic $C(s_{t},a_{t})$, and the target critic $TC(s_{t},a_{t})$.
  • The actor $A(s_{t})$ explores and learns a policy $\pi$ that maps the agent’s observation of the state to the action space $a$. For exploring the optimal policy $\pi^{*}$, the actor selects actions $a_{t}\in a$ based on its current policy, and progressively refines them to maximize the long-term expected rewards. The optimization objective of the agent is expressed as $\pi^{*}(a_{t}|s_{t})=\arg\max_{a_{t}}\mathbb{E}\left[\sum_{t=0}^{\infty}\gamma^{t}r_{t}\right]$, where $\gamma$ is the discount factor that controls the contribution of future rewards. Also, $\mathbb{E}[\cdot]$ denotes the expected value operator.
  • The critic $C(s_{t},a_{t})$ estimates the state-action value, which evaluates the quality of the actions selected by the actor in state $s_{t}$ at time $t$. This value guides the optimization of the actor’s policy towards actions that maximize future expected rewards.
  • To mitigate the instability in the critic arising from frequent updates in the highly dynamic SAGIN environment, $TC(s_{t},a_{t})$ is introduced to compute the Bellman equation-based estimates of future state-action values. When the agent takes an action $a_{t}$ at time $t$ by transmitting traffic over the network, the environment’s feedback is sent to $TC(s_{t},a_{t})$ to estimate the corresponding future state-action value. This feedback includes the immediate reward, $r_{t},$ and the next state of the environment, represented by the updated transmission path conditions. Future state-action values are computed as the sum of the immediate reward received at the current step and the discounted estimate of the next state-action value predicted by the target critic network.
@@ IV Case Study of the SAGIN-enabled MC
In this section, we present a case study demonstrating the effectiveness of the proposed approach for optimizing SAGIN-enabled MC. To this end, we conduct a simulation in a stochastic SAGIN environment comprising ground, aerial, and space platforms. Each episode consists of 50 time-steps, during which an agent selects any non-empty subset of the four candidate links from BS, UAV, HAP, and LEO satellite. We model the line-of-sight (LOS) probability of a link using standard propagation formulations from 3GPP TR 38.901 [12] and ITU-R P.1410/P.618 [13], as well as geometry-based models, as in [14] and [15].
@@
Figure 4: The illustration of the learning behavior and convergence performance of the agent: (a) episode reward, (b) switching rate, (c) average latency, and (d) normalized power consumption.
@@ IV-A Learning Behvaior and Convergence
Fig. 4(a) shows the total reward plotted over training episodes for the SAGIN-enabled MC using the proposed algorithm. Both the actual values and their moving average computed over the last ten samples are presented. Initially, within the first 20 episodes, the agent starts with random actions and poor policies, leading to low rewards due to high latency, low capacity, and inefficient power use. As the agent learns from environmental interactions, its decision-making regarding link selection and resource utilization improves and optimizes. This results in a gradual increase in the cumulative reward, indicating its learning progress.
Fig. 4(b) illustrates the frequency with which the agent switches between available links during each episode in the SAGIN-enabled MC environment using the proposed algorithm. Initial episodes exhibit high or fluctuating switching rates as the agent explores various link combinations, resulting in frequent switching while searching for optimal connections. This behavior reflects exploration. As learning progresses, the agent identifies which links yield stable, high rewards.
Fig. 4(c) shows the latency values and their moving average, as obtained by the proposed algorithm, in the given scenario. The decreasing order of the average values indicates that, as training progresses, the agent learns to minimize latency. Similarly, Fig. 4(d) shows the average transmit power in watts across training episodes. The gradual decline indicates that the RL agent is progressively learning to optimize link selection and resource allocation, thereby reducing overall power consumption. The initial rise and subsequent fall in power reflect the exploration-exploitation dynamics of the proposed algorithm.
@@ IV-B Performance Evaluation
We compare the proposed algorithm against random link selection, round-robin single-link cycling, and greedy single-link selection based on the highest signal-to-noise ratio (SNR). We also compare it against a base station (BS)-only conservative terrestrial baseline.
@@ IV-B1 Capacity
Fig. 5(a) depicts the average capacity achieved by the proposed and baseline schemes. The proposed scheme achieves the highest capacity thanks to its ability to exploit the most favorable combination of ground, aerial, and space links simultaneously. The single-best (based on the highest SNR) scheme ranks second, but it is limited by single-link selection. This restricts diversity and link aggregation gains. In contrast, the random and round-robin approaches suffer from inefficient link usage and frequent suboptimal selections, resulting in significantly lower capacities. The BS-only configuration yields the lowest capacity because it lacks access to the additional spectral and spatial resources of UAV, HAP, and LEO platforms.
@@ IV-B2 Latency
Fig. 5(b) compares the average end-to-end latency. The proposed scheme achieves the lowest latency by intelligently selecting low-delay ground and aerial links while avoiding congested or high-propagation-delay connections. In contrast, the Random and Round-Robin schemes incur delays that are several orders of magnitude higher due to inefficient link utilization and the selection of long-distance or weak channels. The single-best scheme performs moderately well, yet it is suboptimal because it disregards latency-power trade-offs. The BS-only scheme yields stable, albeit higher, latency than the proposed scheme because it lacks multi-link connectivity.
@@ IV-B3 Power Consumption
Fig. 5(c) compares average power consumption. The proposed scheme has a higher average power cost because it uses multiple links. In contrast, the single-best and BS-only schemes consume less power, and the random and round-robin methods consume an intermediate amount of power. The higher power cost of the proposed scheme is due to the activation of multiple simultaneous connections; the baseline schemes use single links. Overall, however, the proposed scheme delivers higher capacity and lower latency with a moderate increase in power consumption. Conversely, the single-best and BS-Only schemes deliver significantly lower capacity, while the random and round-robin methods demonstrate poor latency performance.
@@ Acknowledgment
This work was supported by the National Research Foundation of Korea (NRF) grant funded by the Korean government (MSIT) (No. 2019R1proposed2007037 and 2022R1A4A3033401) and by the MSIT (Ministry of Science and ICT), Korea, under the ITRC (Information Technology Research Center) support program (IITP-2021-0-02046) supervised by the IITP (Institute for Information &\& Communications Technology Planning &\& Evaluation).

### v2

@@ Abstract
Space-air-ground-integrated network (SAGIN)-enabled multiconnectivity (MC) is emerging as a key enabler for next-generation networks, enabling users to simultaneously utilize multiple links across multi-layer non-terrestrial networks (NTN) and multi-radio access technology (multi-RAT) terrestrial networks (TN). However, the heterogeneity of TN and NTN introduces complex architectural challenges that complicate MC implementation. Specifically, the diversity of link types, spanning air-to-air, air-to-space, space-to-space, space-to-ground, and ground-to-ground communications, renders optimal resource allocation highly complex. Recent advancements in agentic reinforcement learning (RL) have shown remarkable effectiveness in optimal decision-making in complex and dynamic environments. In this paper, we review the current developments in SAGIN-enabled MC and outline the key challenges associated with its implementation. We further highlight the transformative potential of learning-driven approaches for resource optimization in a heterogeneous SAGIN environment. To this end, we present a case study on resource allocation optimization enabled by agentic RL for SAGIN-enabled MC involving diverse RATs. Results show that agentic RL can effectively handle complex scenarios and substantially improve latency and capacity with reduced power consumption. Finally, open research problems and future directions are presented to realize efficient SAGIN-enabled MC.
@@ I-A MC in the Unified and Standardized SAGIN Architecture
Nonetheless, the integration of multi-layer NTNs and multi-RAT TNs introduces complex architectures and a plethora of use cases. Consequently, a variety of link types include air-to-air, air-to-space, space-to-space, space-to-ground, and ground-to-ground links. Fig. 1 illustrates a scenario of the heterogeneous environment and architectures enabled by unified SAGIN, where the coexistence of space, air, and ground platforms is considered. As shown in the figure, it encompasses unmanned aerial vehicles (UAVs), high-altitude platforms (HAPs), satellites, and hybrid gateway-base station (GBS) with highly dynamic and complex network conditions, which makes it more challenging to optimize resource allocation and management. For instance, rapidly changing channel conditions, frequent handovers across space, air, and ground domains, and varying latency, bandwidth, and reliability requirements are the factors that contribute to the resource management complexity.
@@
  • The proposed algorithm is validated through simulations and compared with both model- and learning-driven baselines. In addition, we present a comprehensive overview of open challenges and potential research directions.
@@ III Agentic RL for SAGIN-enabled MC
Recently, agentic RL has emerged as a key tool for addressing complex network optimization challenges [11]. In the context of SAGIN-enabled MC, such agents optimize resource allocation under dynamic and complex conditions. The agents continuously learn rapidly changing network states, diverse user requirements, and complex mobility patterns. This enables them to adapt to environmental dynamics and facilitate efficient service delivery. The agents also enable proactive decision-making regarding link activation, handover, traffic steering, and load balancing to predict network conditions. This allows for the intelligent definition of dynamic resource allocation patterns and the optimization of parameters such as transmit power, bandwidth, and carrier selection to maximize spectral and energy efficiency. Thus, agentic RL can enable intelligent, self-optimizing control policies for SAGIN-enabled MC. These policies can orchestrate handovers, link selection, resource allocation, and load balancing across heterogeneous domains while ensuring scalability and adaptability under highly nonlinear dynamics.
@@
Consequently, we have $\{\operatorname{\mathcal{S}},\operatorname{\mathcal{A}},p(s^{\prime}_{t}|s_{t},a_{t}),r_{t}\}$, where $\operatorname{\mathcal{S}}$ denotes the state space of the environment, while $\operatorname{\mathcal{A}}$ represents the action space for making the optimal decision. In addition, $p(s^{\prime}_{t}|s_{t},a_{t})$ corresponds to the probability of transitioning to state $s^{\prime}_{t}\in\operatorname{\mathcal{S}}$ conditioned on taking action $a_{t}\in\operatorname{\mathcal{A}}$ in state $s_{t}\in\operatorname{\mathcal{S}}$. The term $r_{t}$ represents the immediate reward obtained by taking action $a_{t}$, which drives the transition from $s_{t}$ to $s^{\prime}_{t}$. We assume that a UE has access to $N$ available links, indexed by $n\in\{1,...,N\}$, from which a subset can be dynamically assigned to the UE.
@@ III-C Reward
The immediate reward $r_{t}$ is formulated as a penalty imposed on the agent that is proportional to the capacity, latency, and power usage calculated over an episode. More precisely, the reward at each time step is defined as a weighted combination that increases with higher capacity and decreases with higher latency and power consumption. Thus, the reward function maximizes the achievable transmission rate while minimizing system delay. The function implicitly penalizes packet loss, high delays, and exceeding network load constraints, thereby guiding the agent to learn optimal link selection policies. Furthermore, it encourages selecting fewer links to minimize cost.
@@ III-D Agent Architecture
As illustrated in Fig. 3, the proposed learning agent is tailored for SAGIN-enabled MC decision-making, where link availability, traffic dynamics, and resource utilization jointly evolve across heterogeneous space–air–ground layers. The agent architecture comprises three neural networks: the actor $A(s_{t})$, the critic $C(s_{t},a_{t})$, and the target critic $TC(s_{t},a_{t})$.
  • The actor learns a stochastic policy $\pi$ that maps the agent’s observation of the state i.e., channel state, traffic load, resource utilization, connectivity status, and previous action to the action space $a_{t}$ representing link selection and scheduling decisions across available SAGIN links. Specifically, the actor determines which subset of satellite, HAP, UAV, and terrestrial links should be activated at time $t$ to satisfy QoS requirements. The policy is optimized to maximize the expected long-term cumulative reward, $\pi^{*}(a_{t}|s_{t})=\arg\max_{a_{t}}\mathbb{E}\left[\sum_{t=0}^{\infty}\gamma^{t}r_{t}\right]$, where $\gamma$ is the discount factor that controls the contribution of future rewards. Also, $\mathbb{E}[\cdot]$ denotes the expected value operator.
  • The critic estimates the state-action value, quantifying the quality of the actor’s link-selection and scheduling decisions in terms of jointly optimized KPIs, i.e., capacity, latency, and power consumption. This way, the critic guides the actor towards policies that are robust to frequent topology changes and traffic fluctuations in SAGIN.
  • To stabilize learning in the highly dynamic SAGIN environment, a target-critic is introduced to compute the Bellman equation-based estimates of future state-action values. After action $a_{t}$ is taken, i.e., utilization of the selected MC links started, the environment returns the immediate reward $r_{t}$ and the next state $s^{\prime}_{t}$ reflecting updated link conditions and resource availability. The target-critic estimates future state-action values as the sum of the immediate reward and the discounted value of the next state, thereby mitigating oscillations caused by rapidly varying channel and mobility conditions.
@@ IV Case Study of the SAGIN-enabled MC
In this section, we present a case study demonstrating the effectiveness of the proposed approach for optimizing SAGIN-enabled MC. To this end, we conduct a simulation in a stochastic SAGIN environment comprising ground, aerial, and space platforms. The agent and scheduler are deployed at a centralized controller in GBS, where link-selection decisions are made based on channel availability and quality, previous actions, platform load, and resource utilization. Each episode consists of 50 time-steps, during which an agent selects any non-empty subset of the four candidate links from BS, UAV, HAP, and LEO satellite. We model the line-of-sight (LOS) probability of a link using standard propagation formulations from 3GPP TR 38.901 [12] and ITU-R P.1410/P.618 [13], as well as geometry-based models, as in [14] and [15].
@@
Figure 4: The illustration of the learning behavior and convergence performance of the agent: (a) episode reward, (b) switching rate, (c) average latency, and (d) power consumption.
@@ IV-A Learning Behvaior and Convergence
Fig. 4(a) shows the total reward/return plotted over training episodes for the SAGIN-enabled MC. The proposed algorithm converges rapidly compared to DQN and PPO, which is attributable to its stabilizing components that keeps the policy updated in highly dynamic environments. DQN exhibits slower convergence and higher variability due to its off-policy nature, which cannot adapt efficiently to the highly dynamic SAGIN environment. Similarly, the inefficient optimization policy of PPO in the highly dynamic SAGIN environment limits its ability to fully exploit favorable link combinations, resulting in a lower return.
Fig. 4(b) illustrates the links switching rate among the available links during each episode. The proposed algorithm increasingly reduces the switching rate, converging toward near-zero values, as training progress. DQN consistently maintains high switching rate, indicating it struggle with optimal links selection in the highly dynamic SAGIN environment. PPO shows the lowest switching rate. However, given its low return values, this lowest switching rate can be interpreted as locking to fixed and sub-optimal links selection, which results in low return.
In Fig. 4(c), the proposed scheme consistently achieves lower end-to-end latency. DQN performs better than PPO. The conservative updates in PPO prevent it from exploiting low-latency terrestrial and aerial links when conditions are favorable. Similarly, Fig. 4(d) shows the average transmit power in watts. The proposed method maintains lower power usage than both DQN and PPO. DQN utilizes high-power aerial and satellite links during exploration, leading to persistently higher energy expenditure. PPO performs better but still incurs higher power usage due to its slower policy adaptation.
@@ IV-B Performance Evaluation
We compare the proposed algorithm against random link selection, round-robin single-link cycling, greedy single-link selection based on the highest signal-to-noise ratio (SNR), BS-only conservative terrestrial baseline, DQN, and PPO.
@@ IV-B1 Capacity
Fig. 5(a) depicts the average capacity performance. The proposed scheme excels all the competing schemes due to intelligently selecting the optimal set of ground, aerial, and space links simultaneously. The single-best (based on the highest SNR) scheme stands as the second-best performing, however, it is limited by single-link selection. The random and round-robin approaches suffer from inefficient link usage and suboptimal selections, resulting in significantly lower capacities. The single-BS configuration yields the lowest capacity as it lacks access to the additional spectral and spatial resources of UAV, HAP, and LEO platforms. DQN performs better than random and round-robin; however, its performance is limited by its limited action representation and suboptimal exploitation of multiple links. PPO performs lower due to its conservative policy updates and higher sensitivity to delayed rewards in highly heterogeneous SAGIN environments.
@@ IV-B2 Latency
In Fig. 5(b) compares the average end-to-end latency. The proposed scheme achieves the lowest latency by intelligently selecting low-delay ground and aerial links. DQN achieves low latency compared with other schemes. In contrast, PPO experiences high latency. The Random and Round-Robin schemes incur delays that are several orders of magnitude higher due to inefficient link utilization and the selection of long-distance or weak channels. The single-best scheme performs moderately well, yet it is suboptimal as high-SNR links, particularly NTN links, still incur large propagation delays. The single-BS scheme yields stable yet higher latency than the proposed scheme as it lacks multi-link connectivity.
@@ IV-B3 Power Consumption
Fig. 5(c) compares average power consumption. The proposed scheme has a higher average power cost than single-BS because of using multiple links. The single-best and single-BS schemes also consume comparatively low power as they rely on only one active link. The random and round-robin methods consume an intermediate amount of power. DQN shows moderate power usage, while PPO consumes comparatively higher power due to its less stable policy behavior and inefficient link activation.
@@ Acknowledgment
This work was supported by the National Research Foundation of Korea (NRF) grant funded by the Korean government (MSIT) (No. 2019R1proposed2007037 and 2022R1A4A3033401) and by the MSIT (Ministry of Science and ICT), Korea, under the ITRC (Information Technology Research Center) support program (IITP-2021-0-02046) supervised by the IITP (Institute for Information &\& Communications Technology Planning &\& Evaluation). This work was also supported in part by the SNS JU European Union’s Horizon Europe Research and Innovation Program under Grant 101139194 (6GXCEL); in part by the Horizon Europe Program under the MSCA Staff Exchanges under Grant 101086218 (EVOLVE); in part by FWO, Belgium, under Grant for a scientific stay in Flanders (file number V507025N).